From nch@roe.ac.uk Fri Nov  2 16:53:49 2007
Date: Fri, 2 Nov 2007 16:48:05 +0000 (GMT)
From: Nigel Hambly <nch@roe.ac.uk>
To: WFCAM Science Archive Team -- Eckhard Sutorius <etws@roe.ac.uk>,
     Johann Bryant <jb@roe.ac.uk>, Mike Read <mar@roe.ac.uk>,
     Nigel Hambly <nch@roe.ac.uk>, Nicholas Cross <njc@roe.ac.uk>,
     Bob Mann <rgm@roe.ac.uk>, Ross Collins <rsc@roe.ac.uk>
Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence
    <al@roe.ac.uk>, Andy Adamson <a.adamson@jach.hawaii.edu>,
     Brian Walshe <bcw@roe.ac.uk>, Jim Emerson <j.p.emerson@qmul.ac.uk>,
     Malcolm Stewart <jms@roe.ac.uk>, Lorenzo Rimoldini <lgr@roe.ac.uk>,
     Mike Irwin <mike@ast.cam.ac.uk>, Mark Holliman <msh@roe.ac.uk>,
     Peredur Williams <pmw@roe.ac.uk>, Stephen Warren <s.j.warren@ic.ac.uk>
Subject: WFAU VDFS Science Archive weekly project meeting minutes, 2/11/07

Minutes of WFAU VDFS Science Archive meeting:   2nd November 2007
-------------------------------------------------------------------------
-------------------------------------------------------------------------

Present:       NCH, RSC, NJC, MAR, PMW, JB, ETWS, LGR, MSH
Apologies:     BCW, JPE, AL, RGM, JMS

NB: DONM: 10am, THURSDAY 8th November 2007 in the Plate Library


Actions discharged this week:
-----------------------------

ACTION: JB to ngest CU4 non-survey data before running detection table
         shrink on G:\
Discharged.


Actions partly discharged but continuing:
-----------------------------------------

The following from last time partly done but continue:

o) Add in a default row for every detector appearing in every detection
table (for schema consistency when querying merged sources and 
individual detections)
  - ACTION: RSC & NCH
Continues; will get sorted (honestly!) as part of the general
ingest DB schema revamp to facilitate rapid generalised
photometric/astrometric recalibration after DR3.

ACTION: RSC to progress implementation of automated replication of
         the ingest DB file metadata on a public server.
Continues; RSC has located and fettled the appropriate scripts, and
will now test against the restored ingest DB. A similar
solution is proposed for proprietary-lapsed non-survey data
access (see Non-survey Data Release below). Some discussion took
place as to the names of these databases: FlatFiles? WFCAMfiles? ...?


Actions carried forward from 26/10/07 meeting:
----------------------------------------------

ACTION: JB & ETWS to restore SegueDR6 on thutmose at their convenience
         (i.e. no hurry given the present circumstances)
Continues; team agreed to postpone until after DR3 now.

ACTION: JB to set dfg_1 shrinking in the WSA this weekend.
Continues; was not possible to do this as still fettling ingestions
from fixed data.


Specific points and new actions:
--------------------------------

Project management:

NCH & PMW noted the VDMT of the 8th; PMW will assemble the usual
stuff.


WFCAM & VISTA updates:

Congratulations to JAC on achieving the one millionth WFCAM observe
this week.


Comments and issues arising from CASU fortnightly minutes:

No new minutes this week.


Networking:

07B transfers continue over UKLight. Disk copy from old file server 
sneferu to NAS box is clogging things a bit, but when finished
will enable further cross-mount removal.


WSA Operations:

UKIDSS DR3 preparations continue; mostly QC this week.


Hardware

Premises have informed NCH that there will be power downtime in C1
on the 8th of December, subject to confirmation by external parties
(i.e. Scottish Power) and operational needs (i.e. Scuba-2 cooldowns).

ACTION: MAR to put a note on the WSA downtime pages.

PMW asked about the situation as regards retiring the old SuperCOSMOS
webserver cosaxp6. This will be revisited after DR3 and when JB and
MSH have had a chance to think further on whether we go for more
NAS boxes or SAS additions to existing ones.

ACTION: JB & MSH to make a decision as regards the next chunk of
         archive mass-file storage (SAS/NAS...)

NCH noted that the aircon unit is icing up again (maybe due to the
unseasonably mild weather at the moment).

ACTION: JB to manually defrost the aircon unit behind the archive
         servers lest it start peeing water on them from a great height


Software:

NJC reported a bit of wwork on CU6 and galaxy photometry; RSC 
reported a DR3 SVN release branch readiness review.


Survey Data Release:

Updated schedule following on from recent work over the last week:

Weeks What:

  0.2  QC1
  0.0  CU5 (diff images for the GPS - after QC, but in parallel with the
            following and it's very quick anyway)
  1.0  Quality bit flagging
  3.0  CU7 (source merging from scratch again, unfortunately)
  0.0  CU13/14 for DXS/UDS (in parallel with shallow survey CU7s)
  2.0  Final CUs.

MAR noted that QC1 is basically done, but NCH noted that it hasn't
been possible to close out QC fully this week so more work needed.
NCH and MAR need to apply the final cross-semester
deprecations next week. NCH has had some discussions with SJW
about the retiling problem in the LAS; the approach will be 
to use RSC's new CU7 option to create a DR3 frameset table
without source merging in the first instance to check exactly
those frames that should be cleared out from early runs in
retiled areas. NJC reported that UDS folks Omar & Seb have
promised us the DR3 UDS stacks by next Thursday, at which point
NJC will take over and do his usual thing.


Non-survey Data Release:

MAR and JB noted that all current registrations are now set up
and enabled for flat-file access.


Astrogrid deployment:

No news this week.


Miscellaneous:

Nothing else this week